OpenStudy (anonymous):

What is the solution for x in the proportion 2/4 = x/10? choices: 2 4 20 5

OpenStudy (anonymous):

4. you have to make the denominators equal so the lcd is 40 we assume x is one and multiply it by four bc we multiplied the denominator by 4 :)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

no problem :)

Parth (parthkohli):

@rebeccaskell94 better luck next time ;) 4x = 10 * 2 4x = 20 x = 20/4 x = 5

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

meh bad